i need a little help with adding or subtracting unlike rational expressions
2/5w+10-3/2w-4

Answer by fractalier(1829)

If this is what you mean
2 / (5w + 10) - 3 / (2w - 4)
then you have to find a lowest common denominator...we can factor denominators to see it better...we get
2 / 5(w + 2) - 3 / 2(w - 2)
the LCD would then be 10(w + 2)(w - 2)...convert each fraction to that denom...and get
4(w - 2) / 10(w + 2)(w - 2) - 15(w + 2) / 10(w + 2)(w - 2) =
[4(w - 2) - 15(w + 2)] / 10(w + 2)(w - 2) =
(4w - 8 - 15w - 30) / 10(w + 2)(w - 2) =
(-11w - 38) / 10(w + 2)(w - 2)
and we're done